# Descent Methods for Vector Optimization Problems: A Majorization-Minimization Perspective

### v1

@@ 8             $x^{k+1}:=x^{k}+t_{k}d^{k}$
where $C:=conv\{A_{i},i\in[l]\}$ is a base of $K^{*}$. In other words, selecting a transform matrix in (12) is equivalent to selecting a base of $K^{*}$. As a result, the linear convergence rate of Algorithm 7 is sensitive to the choice of $A$. When $K=\mathbb{R}^{m}_{+}$, the steepest descent method [16] fixs $A=I_{m}$, i.e., $C=\Delta_{m}$.
@@ 6.2 Barzilai-Borwein descent method for VOPs with polyhedral cones
In general, for the $e_{k}$ defined as (11) we have $-\bm{\ell}\preceq_{K}e_{k}\preceq_{K}\bm{\ell}$, which can be written as $-A\bm{\ell}\preceq Ae_{k}\preceq A\bm{\ell}$. We denote $\alpha^{k}\in\mathbb{R}^{l}$ as follows:
@@
where $\Delta^{\alpha^{k}}_{l}:=\{c^{*}\in\mathbb{R}^{l}_{++}:\left\langle{c^{*},\alpha^{k}}\right\rangle=1\}$. The subproblem can be reformulated as follows:
@@
The dual problem (DP) is lower dimensional quadratic programming with unit simplex constraint (the vertices of unit simplex constraint are known), then it can be solved by Frank-Wolfe/conditional gradient method efficiently [see, e.g., 36, 7]. However, dual problem of (14) reads as
$\min\limits_{\lambda\in\Delta_{l}}\frac{1}{2}\left\|\frac{(AJF(x^{k}))^{T}\lambda}{\sum\limits_{i\in[l]}\lambda_{i}\alpha^{k}_{i}}\right\|^{2},$
@@
If $K=\mathbb{R}^{m}_{+}$ and $A=I_{m}$, the subproblem (16) reduces to that of the BBDMO [7]. Consequently, transforming (14) into (16) gives a new insight into BBDMO. Comparing the subproblems (12) and (15), we conclude that the differences between the directions of the steepest descent and the Barzilai-Borwein descent depend on the choice of base of $K^{*}$.
@@ 9             $x^{k+1}:=x^{k}+t_{k}d^{k}$
    $A\bm{\mu}\preceq\alpha_{k}\preceq A\bm{\ell}$;
@@
Let $A^{1},A^{2}\in\mathcal{A}$, $d^{k}_{1},t_{k}^{1}$ and $d^{k}_{2},t_{k}^{2}$ be the descent directions and stepsize generated by Algorithm 8 with $A^{1}$ and $A^{2}$, respectively. If $\alpha_{\min}<\alpha_{i}^{k,1},\alpha_{i}^{k,2}<\alpha_{\max}$, $i\in[l]$, we have $d^{k}_{1}=d^{k}_{2}$ and $t_{k}^{1}=t_{k}^{2}$.
@@ 8 Conclusions
In future work, it is worth analyzing proximal gradient methods and second-order methods for VOPs from a majorization-minimization perspective. Furthermore, solution methods for VOPs with non-polyhedral cones are also worth investigating.
@@ Acknowledgement
This work was supported by the Major Program of the National Natural Science Foundation of China [grant numbers 11991020, 11991024]; the National Natural Science Foundation of China [grant numbers 11971084, 12171060]; and the Natural Science Foundation of Chongqing [grant number cstc2019jcyj-zdxmX0016].

### v2

@@ 6 Descent methods for VOPs with polyhedral cones
Notably, in practice, the polyhedral cone $K$ is often defined by a specific transform matrix $A$. This raises a crucial question: does the transform matrix $A$ affect the performance of descent methods for VOPs with polyhedral cones?
@@ 8             $x^{k+1}:=x^{k}+t_{k}d^{k}$
where $C:=conv\{A_{i},i\in[l]\}$ is a base of $K^{*}$. In other words, selecting a transform matrix in (16) is equivalent to selecting a base of $K^{*}$. Consequently, the linear convergence rate of Algorithm 7 is sensitive to the choice of $A$. When $K=\mathbb{R}^{m}_{+}$, the steepest descent method [17] fixs $A=I_{m}$, i.e., $C=\Delta_{m}$.
@@ 6.2 Barzilai-Borwein descent method for VOPs with polyhedral cones
In general, for the $e_{k}$ defined as (15) we have $e_{k}\preceq_{K}\bm{\ell}$, which can be written as $Ae_{k}\preceq A\bm{\ell}$. We denote $\alpha^{k}\in\mathbb{R}^{l}_{++}$ as follows:
@@
where $\Delta^{\alpha^{k}}_{l}:=\{c^{*}\in\mathbb{R}^{l}_{+}:\left\langle{c^{*},\alpha^{k}}\right\rangle=1\}$. The subproblem can be reformulated as follows:
@@
The dual problem (DP) is lower dimensional quadratic programming with unit simplex constraint (the vertices of unit simplex constraint are known), then it can be solved by Frank-Wolfe/conditional gradient method efficiently [see, e.g., 41, 7]. However, dual problem of (18) reads as
$\min\limits_{\lambda\in\Delta_{l}}\frac{1}{2}\frac{\left\|(AJF(x^{k}))^{T}\lambda\right\|^{2}}{\sum\limits_{i\in[l]}\lambda_{i}\alpha^{k}_{i}},$
@@
If $K=\mathbb{R}^{m}_{+}$ and $A=I_{m}$, the subproblem (20) reduces to that of the BBDMO [7]. Consequently, transforming (18) into (20) gives a new insight into BBDMO. Comparing the subproblems (16) and (19), we conclude that the differences between the directions of the steepest descent and the Barzilai-Borwein descent lie in the choice of base of $K^{*}$.
@@ 9             $x^{k+1}:=x^{k}+t_{k}d^{k}$
    $A\bm{\mu}\preceq\alpha^{k}\preceq A\bm{\ell}$;
@@
(Affine Invariance) Let $A^{1},A^{2}\in\mathcal{A}$, $d^{k}_{1},t_{k}^{1}$ and $d^{k}_{2},t_{k}^{2}$ be the descent directions and stepsize generated by Algorithm 8 with $A^{1}$ and $A^{2}$, respectively. If $\alpha_{\min}<\alpha_{i}^{k,1},\alpha_{i}^{k,2}<\alpha_{\max}$, $i\in[l]$, we have $d^{k}_{1}=d^{k}_{2}$ and $t_{k}^{1}=t_{k}^{2}$.
@@
7 Second-order methods from a majorization-minimization perspective
Let $K=\mathbb{R}^{m}_{+}$, Fliege et al. [16] proposed a Newton method for MOPs, the Newton direction is defined as the optimal solution of
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta_{m}}\left\langle{\lambda,JF(x^{k})d+\frac{1}{2}d^{T}D^{2}F(x^{k})d}\right\rangle,$ (24)
where $d^{T}D^{2}F(x^{k})d:=(d^{T}\nabla^{2}F_{1}(x^{k})d,...,d^{T}\nabla^{2}F_{m}(x^{k})d)$. The objective-wise quadratic approximation serves as good surrogate in Newton method but makes the subproblem challenging to solve. From a majorization-minimization perspective, the surrogate should strike a balance between better approximation and lower computational cost. In the remainder of this section, we develop an efficient second-order method. Interestingly, the proposed method coincides with the preconditioned Barzilai-Borwein method for MOPs [9].
Firstly, we note that the challenging nature of the subproblem (24) arises primarily from selecting $\Delta_{m}$ as the base for +$\mathbb{R}^{m}_{+}$. By instead choosing a general base in the subproblem:
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta^{\alpha^{k}}_{m}}\left\langle{\lambda,JF(x^{k})d+\frac{1}{2}d^{T}D^{2}F(x^{k})d}\right\rangle,$ (25)
where $\Delta^{\alpha^{k}}_{m}:=\{c^{*}\in\mathbb{R}^{m}_{+}:\left\langle{c^{*},\alpha^{k}}\right\rangle=1\}$, it can be reformulated as follows:
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta_{m}}\sum\limits_{i\in[m]}\lambda_{i}\frac{\left\langle{\nabla F_{i}(x^{k}),d}\right\rangle}{\alpha^{k}_{i}}+\frac{1}{2}\sum\limits_{i\in[m]}\lambda_{i}\frac{\left\langle{d,\nabla^{2}F_{i}(x^{k})d}\right\rangle}{\alpha^{k}_{i}}.$ (26)
This raises a direct question: how to select an appropriate $\alpha^{k}$? Intuitively, if
$\frac{\left\langle{d^{k},\nabla^{2}F_{i}(x^{k})d^{k}}\right\rangle}{\alpha^{k}_{i}}=\sum\limits_{i\in[m]}\lambda^{k}_{i}\frac{\left\langle{d^{k},\nabla^{2}F_{i}(x^{k})d^{k}}\right\rangle}{\alpha^{k}_{i}}$
for $i\in[m]$, by setting $B_{k}\approx\sum_{i\in[m]}\lambda^{k}_{i}{\nabla^{2}F_{i}(x^{k})}/{\alpha^{k}_{i}}$, we reformulate (26) into the following subproblem:
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta_{m}}\sum\limits_{i\in[m]}\lambda_{i}\frac{\left\langle{\nabla F_{i}(x^{k}),d}\right\rangle}{\alpha^{k}_{i}}+\frac{1}{2}\left\lVert{d}\right\rVert^{2}_{B_{k}},$ (27)
where $\alpha^{k}_{i}\approx\left\langle{d^{k},\nabla^{2}F_{i}(x^{k})d^{k}}\right\rangle/\left\lVert{d^{k}}\right\rVert^{2}_{B_{k}}$ for $i\in[m]$. Since $d^{k}$ is unknown before solving the subproblem, we estimate it with $d^{k-1}$ and set
$\alpha^{k}_{i}=\frac{\left\langle{\nabla F_{i}(x^{k})-\nabla F_{i}(x^{k-1}),x^{k}-x^{k-1}}\right\rangle}{\left\lVert{x^{k}-x^{k-1}}\right\rVert^{2}_{B_{k}}}$
for $i\in[m]$. Unfortunately, it is important to note that $\lambda^{k}$ is unavailable prior to computing the subproblem, and $\alpha^{k}$ is determined based on $B_{k}$. As an alternative, we can approximate $B_{k}$ using $\sum_{i\in[m]}\lambda^{k-1}_{i}{\nabla^{2}F_{i}(x^{k})}/{\alpha^{k-1}_{i}}$ instead of $\sum_{i\in[m]}\lambda^{k}_{i}{\nabla^{2}F_{i}(x^{k})}/{\alpha^{k}_{i}}$. Consequently, by updating $B_{k}$ via the BFGS method, the subproblem (27) coincides with that of preconditioned Barzilai-Borwein method for MOPs, which can be efficiently solved using the Frank-Wolfe method [9].
Remark 7.1.
Chen et al. [9] provided a detailed explanation of why the combination of the Barzilai-Borwein (BB) method and the BFGS method is essential in the multi-objective optimization context, particularly from a preconditioning perspective. Specifically, these two approaches function as complementary preconditioners for the search direction. While their effects tend to overlap in single-objective optimization, in the multi-objective setting, they fulfill distinct roles and demonstrate improved performance when applied together. In this section, we also highlight the indispensability of the Barzilai-Borwein rule in the quasi-Newton method from a majorization-minimization perspective.
@@ 9 Conclusions
From a majoriztion-minimization perspective, we also rediscover the preconditioned Barzilai-Borwein method for MOPs. This highlights the versatility of the majorization-minimization principle as a powerful framework for designing novel algorithms in vector optimization. In future work, it is worth analyzing proximal gradient methods and high-order methods for VOPs within the majorization-minimization framework. Additionally, exploring solution methods for VOPs with non-polyhedral cones presents an intriguing avenue for further research.
@@ Acknowledgement
This work was funded by the Major Program of the National Natural Science Foundation of China [grant numbers 11991020, 11991024]; the Key Program of the National Natural Science Foundation of China [grant number 12431010]; the General Program of the National Natural Science Foundation of China [grant number 12171060]; NSFC-RGC (Hong Kong) Joint Research Program [grant number 12261160365].